# Rethinking High-speed Image Reconstruction Framework with Spike Camera

### v1

@@ Abstract
* Corresponding authors. Copyright © 2025, Association for the Advancement of Artificial Intelligence (www.aaai.org). All rights reserved.
@@ Methods
Regarding the architecture of the Recon-Net, since our method does not have ground truth as a strong constraint, we design a Lightweight Reconstruction Network (LRN) composed of several convolutional layers as in (Chen et al. 2024) to accomplish this task, which is efficient and sufficient as demonstrated by our further experiments. While the spike stream remains sparse under low-light conditions, we employ the voxel technique widely used in previous event-based vision tasks (Song, Huang, and Bajaj 2022) to squeeze the input length of the spike sequence.
@@
We conduct the comparative experiment based on the Spike-Zoo1 repository, while the overall experimental framework is built on the PyTorch platform and trained based on an NVIDIA 4090 GPU. We employ the Adam optimizer with a learning rate of 4e-5 on the U-CALTECH dataset and 1e-4 on the U-CIFAR dataset. The training process consists of 25 total epochs: 5 epochs for coarse reconstruction, 1 epoch for prompt optimization, and 19 epochs for fine reconstruction. We employ the voxelization technique to convert the spike stream from an initial length of 200 to 50.
$1$ https://github.com/chenkang455/Spike-Zoo
@@
Table 3: Ablation on the network architecture.
@@ Acknowledgments
We sincerely appreciate Yuyan Chen (HUST) for her valuable suggestions and for polishing the figures. This work was supported by the National Natural Science Foundation of China (62176003, 62088102, 62306015), the China Postdoctoral Science Foundation (2023T160015), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Beijing Nova Program (20230484362).

### v2

@@ Methods
Regarding the architecture of the Recon-Net, since our method does not have ground truth as a strong constraint, we design a Lightweight Reconstruction Network (LRN) composed of several convolutional layers as in (Chen et al. 2024) to accomplish this task, which is efficient and sufficient as demonstrated by our further experiments. While the spike stream remains sparse under low-light conditions, we employ the voxel technique widely used in previous event-based vision tasks (Song, Huang, and Bajaj 2022; Chen and Yu 2024) to squeeze the input length of the spike sequence.
@@
We conduct the comparative experiment based on the Spike-Zoo (Chen et al. 2025) repository, while the overall experimental framework is built on the PyTorch platform and trained based on an NVIDIA 4090 GPU. We employ the Adam optimizer with a learning rate of 4e-5 on the U-CALTECH dataset and 1e-4 on the U-CIFAR dataset. The training process consists of 25 total epochs: 5 epochs for coarse reconstruction, 1 epoch for prompt optimization, and 19 epochs for fine reconstruction. We employ the voxelization technique to convert the spike stream from an initial length of 200 to 50.
@@
Table 2: Ablation on the contribution of each component.
@@ Acknowledgments
We sincerely appreciate Yuyan Chen (HUST) for her valuable suggestions and for polishing the figures. This work was supported by the National Natural Science Foundation of China (62422601, 62176003, 62088102, 62306015), the China Postdoctoral Science Foundation (2023T160015), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), and the Beijing Nova Program (20230484362).